A wellknown mineral flourite is chemicallycalciumfloride.itis knownthat in one unit cell of this mineral there are 4ca2+ ions and 8fnegative ions and that ca2+ions are arrangedin a fcc lattice.The f_ions fill all the tetrahedral holesin fcc latticeof ca2+ions.The edge length if unitcell is 5.46×10^-8 cm in length.Thedensity of the.solid is 3.18gcm-3 .use thus information to calculate Avogadro;s number.(molarmass of caF2=78.8)
1
Expert's answer
2017-03-20T09:59:06-0400
p=m/V

4 molecule CaF 2 per lattice.

3.18=4*78.8/(N a *(5.46*10^[-8] )^3 )

N a =6*10^23